How much energy is emitted out when a mole of electron falls in 1g atom of H from 2 to 1 orbit?(Rh=109678 cm-1)

Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Use 1/l=RZ2(1/n1^2-1/n^2)

E=hc/l

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vikas askiitian expert
13 years ago

E = 13.6 (1-1/n2)       

this is the energy librated when 1 electron falls from n=n to n=1

 

for 1 mol of electrons or  Na number of electrons energy will be

ET = 13.6 Na(1-1/n2)

n = 2      (given)

Na = 6 * 1023

 

ET = 13.6*(3/4)*6*1023

ET = 6.12*1024 ev

1 ev = 1.6*10-19 J so

ET = 9.8*105 J or 980  KJ